OpenStudy (anonymous):

-8x-2 = 6x +5 solve for x please

OpenStudy (diyadiya):

First subtract 6x from both sides :D can you do that?

OpenStudy (anonymous):

-8x-6x-2=5 -14x-2=5 -14x=5+2 -14x=7 then i assume you can do the rest

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-2??

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

oh.....my bad

OpenStudy (anonymous):

so is it 2........-3.....or 1/2

OpenStudy (anonymous):

5+2 = 7 @Hally

OpenStudy (anonymous):

im fluttering confused yall

OpenStudy (anonymous):

its -1/2

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

hahahah i tried to type sh**# and it came out as pellet.........you cant swear

OpenStudy (anonymous):

@Hally also -8x-6x = -14x not -2

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

oh my bad... :) hehe sorry.

OpenStudy (anonymous):

its -1/2 :) tim is right :)

OpenStudy (anonymous):

k thanks pellet feather hoe flutter

OpenStudy (anonymous):

sorry about that xDD

OpenStudy (anonymous):

Hi @barbiejazzybabii - that type of language (regardless of the fact that the filter altered it) is NOT acceptable on OpenStudy, understood?

OpenStudy (anonymous):

Well, that was a simple decision. I'm not deleting this question due to the work that was put in by other users. Also, I am particularly in favor of the approach that @Diyadiya *attempted* to use in leading this asker to the solution, as opposed to simply providing her the answer she wanted. Thanks everyone!